OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

each next term is 1/3 of its previous term. so, after 24 will be 24/3 = 8. after 8 will be 8/3. after 8/3 will be 8/9. rule can be n = 1/3*(n-1) where n is the number of term. so that for 4th term, we need have 3rd term & divide it by 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np, you have more questions??

OpenStudy (anonymous):

yes just posted one